# natural language processing; greenwashing detection; corporate climate-related greenwashing; climate pretrained models; climate topic detection; green claim detection; climate QA; climate-related deceptive techniques; enviromental performance prediction copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Natural language processing

### v1

@@ 4.8. Question Answering
The main application of Question-Answering (QA) systems is to analyze complex documents, such as corporate climate reports, that can be particularly useful for greenwashing detection. Modern LLMs achieve strong results in QA due to instruction-following fine-tuning, which improves adherence to specific instructions, and Retrieval-Augmented Generation (RAG), which bases its answers on retrieved documents. However, retrieval systems remain a performance bottleneck (Maekawa et al., 2024). Furthermore, answer generation is challenging to evaluate (Sai et al., 2022), often requiring human feedback or advanced model-based assessments.
4.9. Classification of Deceptive Techniques
In analyzing climate-related discourse, it can be useful to identify rhetorical strategies that might obscure or misrepresent an entity’s stance. For example, expressing support for climate remediation policies on one side (Morio and Manning, 2023), and promoting arguments downplaying the urgency of climate change on the other (Coan et al., 2021) could signal a lack of authenticity. Another example could be promoting a product through misleading rhetoric, by, for example, claiming that a product is better because it is natural, which would be an “appeal to nature” – a fallacious argument (Vaid et al., 2022; Jain et al., 2023). Detecting these deceptive techniques could help identify misleading communications.
@@ 4.9. Classification of Deceptive Techniques
LogicClimate (Jin et al., 2022) texts from climatefeedback.org Faulty Generalization, Ad Hominem, Ad Populum, False Causality, … Classifies fallacies (Multi-label)
Neutralization Techniques (Bhatia et al., 2021) paragraphs from other previous works on climate-change Denial of Responsibility, Denial of Injury, Denial of Victim, Condemnation of the Condemner, … Classifies neutralization techniques
Contrarian Claims (Coan et al., 2021) paragraphs from conservative think tank No Claim, Global Warming is not happening, Climate Solutions won’t work, Climate impacts are not bad, … Classifies arguments into super/sub-categories of climate science denier’s arguments
Table 9. Summary of datasets, labels, scopes, and sources for tasks related to deceptive techniques in climate-related context.
@@
Unsupervised Solutions
Oppong-Tawiah and Webster (2023) frame greenwashing as fake news. They propose to tackle this identification through the form perspective. They developed a profile-deviation-based method to detect greenwashing in corporate tweets by comparing linguistic cues (e.g., quantity, specificity, complexity, diversity, hedging, affect, and vividness) to theoretically ideal profiles of truthful and deceptive communication. They compute a greenwashing score as Euclidean distances.
@@
An organization seeking to justify limited action might adopt a rhetoric that downplays the urgency of global warming or dismisses the impact of negligent behavior. These kinds of climate contrarian arguments mostly fall into a finite set of categories (e.g. “Solutions will not work”, or “Human influence is not demonstrated”). Therefore, Coan et al. (2021) constructed a taxonomy of such contrarian claims and published a large dataset annotated with such claims. Similarly, Bhatia et al. (2021) proposed a dataset to classify neutralization techniques, i.e., rationalizations that individuals use to justify deviant or unethical behavior (e.g. Denial of the victim). Finally, Jin et al. (2022) studied fallacious arguments and how they apply to climate change more broadly. They constructed LogicClimate, a climate-specific dataset of sentences annotated with fallacies, using articles from the Climate Feedback website.
Supervised Solutions
Fine-tuned Transformer architectures and classical approaches have been evaluated on each dataset (Jin et al., 2022; Bhatia et al., 2021; Coan et al., 2021). Jin et al. (2022) experimented with an ELECTRA model, fine-tuned to use the structure of the argument. The performance of fine-tuned approaches ranges from 58.77% to 79%. The models struggle with fallacy detection in LogicClimate, but perform well with generalist fallacy detection, contrarian claims, and neutralization techniques. Thulke et al. (2024) experimented with multiple zero-shot approaches using multiple LLM on the binary classification of contrarian claims using Coan et al. (2021)’s dataset. Their climateGPT-70B and LLama-2-Chat-70B models both reached an F1-score of 72.5% (see Table 23 in the Appendix for details).
@@
The low performance of models on tasks such as fallacy detection and neutralization classification indicates the inherent complexity of these tasks. Fallacy detection, for one, is known to be inherently subjective (Helwe et al., 2024). The detection of neutralization techniques, too, appears to be subjective, as even human annotators achieve only a moderate performance level of 70% F1-score. Furthermore, both LogicClimate (Jin et al., 2022) and the neutralization dataset (Bhatia et al., 2021) are very small in size. To address these limitations, future research could explore several directions, including increasing the size of the datasets, defining the labels more precisely, or permitting multiple correct annotations (Helwe et al., 2024) to acknowledge viable disagreement. Additionally, the analyses on rhetorical techniques could be combined with other analyses (such as stance detection) to search for communication patterns.
4.10. Environmental Performance Prediction
@@
Unsupervised Solutions
Jain et al. (2023) proposed to predict the Scope 3 emissions from the list of financial transactions of the company. They first performed classification of all transaction descriptions to map them to their emission factors (emission per \$ spent), and then compute the emissions of each transaction. For ESG score prediction, Bronzini et al. (2023) proposes to use LLMs as few-shot learners to extract triples of ESG category, action, and company from sustainable reports, and to construct a graph representation with few-shot learning. They demonstrated the triplet generation using Alpaca, WizardLM, Flan-T5, and ChatGPT on a few examples. They could analyze disclosure choices and company similarities using the constructed graph. More importantly, they also used the graph to interpret ESG scores through the interpretability analysis of OLS predictions, effectively predicting the ESG score.
@@
Mehra et al. (2022) proposed to predict the changes in the ESG score instead of the actual value. They constructed their dataset by extracting the three sentences most relevant to the environment from financial reports and associated them with the E score [^5] ESG scores are usually aggregated scores along multiple dimensions. In this study, they are focusing on the ”Environment” part of the ESG score.’s change and direction of change. Instead of relying on scores, Clarkson et al. (2020) proposed to focus on good/bad CSR (corporate social responsibility) performers. Their approach evaluates CSR performance based on linguistic style rather than content, aiming to identify whether language patterns alone influence the perception of CSR quality. Sharma et al. (2024) introduced a multi-modal dataset of the tweets of companies to predict a company’s ESG unmanaged risk. Finally, focusing on products and not companies, Lin et al. (2023b) introduced a dataset of online product descriptions and reviews used to predict the Finch Score. All these scores can be used to identify companies and products that are actually environmentally friendly, helping users distinguish actual sustainability from greenwashing. However, they can also be used to find discrepancies between the communicated green-ness of a product or company (e.g. the percentage of climate-related texts in reports) and the actual green-ness (e.g. ESG score or quantity of emissions), which might indicate cases of greenwashing.
Supervised Solutions
Mehra et al. (2022) evaluated fine-tuned Transformers on ESG change prediction. Clarkson et al. (2020) experimented with hand-chosen features plugged into random forest and SVM on CSR performers prediction. Lin et al. (2023b) also experimented with traditional approaches (e.g. gradient boosting) and custom architectures based on Transformer models on Finch score prediction. Sharma et al. (2024) introduced baseline models, which are simple image and text embedding models (e.g., CLIP or sentence-BERT) with a classification head on ESG score prediction. A detailed list of models is reported in Table LABEL:tab:models in the Appendix. Bronzini et al. (2023); Clarkson et al. (2020); Sharma et al. (2024) did not report baselines, which makes the study difficult to analyze. (Mehra et al., 2022) reported good accuracy, demonstrating that a few sentences hold a significant amount of information to predict ESG score changes. However, there is definitely room for improvement, indicating that the information selected is not sufficient. Lin et al. (2023a) reported the performance of the baseline (average score) reaching already mean squared error (MSE) of 11.7% already quite low. Their Transformer-based approach reached MSE of 7.4% improving slightly on classical approaches (Gradient Boosting reaching MSE of 8.2%). Detailed performances can be found in Table 24 in the Appendix.
@@
We expected the prediction of ESG and CSR metrics to be difficult and require an extensive understanding of a company. However, Bronzini et al. (2023); Lin et al. (2023b); Mehra et al. (2022) and Clarkson et al. (2020) showed that it is possible to build strong predictors relying only on textual elements. This can be explained because analysts reward transparency (Bronzini et al., 2023), so the quantity and complexity of disclosure have strong predictive power (Clarkson et al., 2020). Although existing studies have examined specific dimensions, future research should investigate the interaction between form (Clarkson et al., 2020) and content (Bronzini et al., 2023), as this could help uncover inconsistencies that may indicate greenwashing practices.
5. Greenwashing Detection
@@
There is a large body of work on disclosure (as described in Section 4.3.1). Quantifying climate-related disclosure can highlight greenhushing and selective disclosure. Bingler et al. (2022) identify a lack of disclosure in the Strategy and Metrics&Targets categories, which they describe as cherry-picking. They also highlight that merely announcing support for the TCFD does not lead to an increase in disclosure – on the contrary, it is a practice called cheap talk. Similarly, Auzepy et al. (2023) conducted a more fine-grained analysis of TCFD-related disclosure in the banking industry. They also highlighted large differences across TCFD categories. In particular, they found a low disclosure rate on the fossil-fuel industry-related topic despite large investments in that sector.
@@
Ding et al. (2023) and Chou et al. (2023) found a correlation between climate-related disclosure and carbon emission: companies that emit more tend to disclose more climate-related information. Bingler et al. (2023) proposed a Cheap talk index based on claims specificity. They reached the same conclusion that larger emitters tend to disclose more, but also that negative news coverage is correlated with cheap talk. Polignano et al. (2022) showed that while disclosing more, reports mostly focus on positive disclosure. Kang and Kim (2022) proposed a sentiment ratio metric that highlights overly positive corporate reports. They identified periods of overly positive communications that followed negative environmental controversies; in other words, companies that tried to rebuild their image after a controversy. Gorovaia and Makrominas (2025) analyzed CSR reports of environmental violators and companies with clean records. They studied an ensemble of variables: quantification of environmental content, readability score, and sentiment analysis. They found that violators publish longer, more positive, and less readable CSR reports with more environment-related content. This suggests that companies use climate disclosure as a tool to mitigate controversies and/or to improve their image. Zhao et al. (2023) propose to use the discrepancies between internal disclosure and social media perception of a company to identify potential greenwashing.
@@
Clarkson et al. (2020) found that companies that use a more complex language in their CSR disclosure are associated with a higher CSR rating. Schimanski et al. (2023b) concluded that a higher quantity of ESG communication is associated with higher ESG rating. Rouen et al. (2023) also highlighted the relationship between disclosure quantity and complexity with the ESG score. This might indicate that the linguistic style is a good predictor of the substance of the discourse, or that analysts are rewarding companies that communicate on ESG-related issues.
@@
Lee and Raschke (2023) proposed a greenwashing index based on the difference between the ESG score and the ESG score, weighted by the quantity of communication on each topic (E, S and G). Sharma et al. (2024) define two types of risks: managed risk (the company is addressing it), and unmanaged risk (the company is not currently addressing it but it could). Based on these, they propose an ESG unmanaged risk score as a measure for environmental performance, which can then be used as a signal for identifying greenwashing.
@@
Morio and Manning (2023) proposed to evaluate a company’s stance towards climate change mitigation policies. They hypothesize that companies with a mixed stance might be engaging in greenwashing. The hypothesis was not tested and remains unconfirmed; still, such methodology helps understand the narratives of companies around climate change. Coan et al. (2021) were able to analyze the type of contrarian claims used by conservative think-tank (CTTs) websites and contrarian blogs. They showed a shift over time from climate change denying arguments to opposing climate change mitigation policies. Despite this shift, they found that the primary donors of these think tanks continue to support organizations promoting a climate-denier narrative aimed at discrediting scientific evidence and scientists. Discrepancies between the stance extracted from official communications (Morio and Manning, 2023) on the one hand, and the stance of third-party media organizations such as think tanks founded by the company (Coan et al., 2021) on the other, might indicate greenwashing.
@@
Oppong-Tawiah and Webster (2023) we propose to identify greenwashing using the linguistic profile of tweets. They estimate the deceptiveness of the text using a keyword-based approach and linguistic indicators (e.g. word quantity, sentence quantity). They found a correlation between greenwashing and lower financial performances. A scoring system based solely on linguistic elements has significant limitations. For example, the following tweet gets a high greenwashing score: “Read about #[company’s] commitment to a #lowcarbon future http://[company website]”. However, it is merely an invitation to read a Web page.
@@
Several indicators for greenwashing have been proposed: overly positive sentiment (Kang and Kim, 2022), the use of specific linguistic cues (Oppong-Tawiah and Webster, 2023), discrepancies between ESG scores and corporate disclosures (Lee and Raschke, 2023; Sharma et al., 2024), ambiguous stances (Morio and Manning, 2023), and inconsistencies between social media perception and official disclosures (Zhao et al., 2023). While these approaches laid the theoretical groundwork for understanding indicators of greenwashing, they have a significant limitation: they are not empirically evaluated (Bernini et al., 2023). Only a few studies, such as that by Kang and Kim (2022), have made attempts to validate their signals against actual cases, but even these efforts have not resulted in a comprehensive dataset of greenwashing examples. This gap between theory and practice highlights the necessity of developing datasets containing real-world instances of greenwashing. Without empirical evaluation, the prediction that a company engages in greenwashing is unfounded at best, and misleading or even defamatory at worst. However, building such a dataset comes with many challenges. First, one would have to identify suitable sources of documents that are openly accessible and likely to contain greenwashing. Second, one would have to overcome the inconspicuous and subjective nature of greenwashing itself. Finally, the publication of any such dataset exposes the authors to charges of defamation by the companies they accuse of greenwashing.
6. Open Challenges in Greenwashing Detection
We discuss open challenges in the evaluation (Section 6.1), model robustness (Section 6.2), and data sources (Section 6.3).
6.1. Evaluation Methodology

### v2

@@ 4.8 Question Answering
The main application of Question-Answering (QA) systems is to analyze complex documents, such as corporate climate reports, that can be particularly useful for greenwashing detection. Modern LLMs achieve strong results in QA due to instruction-following fine-tuning, which improves adherence to specific instructions, and Retrieval-Augmented Generation (RAG), which bases its answers on retrieved documents. However, retrieval systems remain a performance bottleneck [maekawa-etal-2024-retrieval]. Furthermore, answer generation is challenging to evaluate [survey_nlg_eval], often requiring human feedback or advanced model-based assessments.
4.9 Classification of Deceptive Techniques
In analyzing climate-related discourse, it can be useful to identify rhetorical strategies that might obscure or misrepresent an entity’s stance. For example, expressing support for climate remediation policies on one side morio2023an, and promoting arguments downplaying the urgency of climate change on the other coanComputerassistedClassificationContrarian2021 could signal a lack of authenticity. Another example could be promoting a product through misleading rhetoric, by, for example, claiming that a product is better because it is natural, which would be an “appeal to nature” – a fallacious argument vaid-etal-2022-towards, jain_supply_2023. Detecting these deceptive techniques could help identify misleading communications.
Table 11: Summary of datasets, labels, scopes, and sources for tasks related to deceptive techniques in climate-related context.
@@ 4.9 Classification of Deceptive Techniques
LogicClimate [jin-etal-2022-logical] texts from climatefeedback.org Faulty Generalization, Ad Hominem, Ad Populum, False Causality, … Classifies fallacies (Multi-label)
Neutralization Techniques [bhatia_automatic_2021-1] paragraphs from other previous works on climate-change Denial of Responsibility, Denial of Injury, Denial of Victim, Condemnation of the Condemner, … Classifies neutralization techniques
Contrarian Claims [coanComputerassistedClassificationContrarian2021] paragraphs from conservative think tank No Claim, Global Warming is not happening, Climate Solutions won’t work, Climate impacts are not bad, … Classifies arguments into super/ sub-categories of climate science denier arguments
@@
Annotation-free
divinus_oppong-tawiah_corporate_2023 frame greenwashing as fake news. They propose to tackle this identification through the form perspective. They developed a profile-deviation-based method to detect greenwashing in corporate tweets by comparing linguistic cues (e.g., quantity, specificity, complexity, diversity, hedging, affect, and vividness) to theoretically ideal profiles of truthful and deceptive communication. They compute a greenwashing score as Euclidean distances.
@@
An organization seeking to justify limited action might adopt a rhetoric that downplays the urgency of global warming or dismisses the impact of negligent behavior. These kinds of climate contrarian arguments mostly fall into a finite set of categories (e.g. “Solutions will not work”, or “Human influence is not demonstrated”). Therefore, coanComputerassistedClassificationContrarian2021 constructed a taxonomy of such contrarian claims and published a large dataset annotated with such claims. Similarly, bhatia_automatic_2021-1 proposed a dataset to classify neutralization techniques, i.e., rationalizations that individuals use to justify deviant or unethical behavior (e.g. Denial of the victim). Finally, jin-etal-2022-logical studied fallacious arguments and how they apply to climate change more broadly. They constructed LogicClimate, a climate-specific dataset of sentences annotated with fallacies, using articles from the Climate Feedback website.
Models
Fine-tuned Transformer architectures and classical approaches have been evaluated on each dataset [jin-etal-2022-logical, bhatia_automatic_2021-1, coanComputerassistedClassificationContrarian2021]. jin-etal-2022-logical experimented with an ELECTRA model, fine-tuned to use the structure of the argument. The performance of fine-tuned approaches ranges from 58.77% to 79%. The models struggle with fallacy detection in LogicClimate, but perform well with generalist fallacy detection, contrarian claims, and neutralization techniques. thulke2024climategpt experimented with multiple zero-shot approaches using multiple LLM on the binary classification of contrarian claims using coanComputerassistedClassificationContrarian2021’s dataset [^9] Their climateGPT-70B and LLama-2-Chat-70B models both reached an F1-score of 72.5% (see Table in the Appendix for details)..
@@
The low performance of models on tasks such as fallacy detection and neutralization classification indicates the inherent complexity of these tasks. Fallacy detection, for one, is known to be inherently subjective [helwe-etal-2024-mafalda]. The detection of neutralization techniques, too, appears to be subjective, as even human annotators achieve only a moderate performance level of 70% F1-score. Furthermore, both LogicClimate [jin-etal-2022-logical] and the neutralization dataset [bhatia_automatic_2021-1] are very small in size. To address these limitations, future research could explore several directions, including increasing the size of the datasets, defining the labels more precisely, or permitting multiple correct annotations [helwe-etal-2024-mafalda] to acknowledge viable disagreement. Additionally, the analyses on rhetorical techniques could be combined with other analyses (such as stance detection) to search for communication patterns.
4.10 Environmental Performance Prediction
@@
Annotation-free
jain_supply_2023 proposed to predict the Scope 3 emissions from the list of financial transactions of the company. They first performed classification of all transaction descriptions to map them to their emission factors (emission per \$ spent), and then compute the emissions of each transaction. For ESG score prediction, bronzini_glitter_2023 proposes to use LLMs as few-shot learners to extract triples of ESG category, action, and company from sustainable reports, and to construct a graph representation with few-shot learning. They demonstrated the triplet generation using Alpaca, WizardLM, Flan-T5, and ChatGPT on a few examples. They could analyze disclosure choices and company similarities using the constructed graph. More importantly, they also used the graph to interpret ESG scores through the interpretability analysis of OLS predictions, effectively predicting the ESG score.
@@
Mehra_2022 proposed to predict the changes in the ESG score instead of the actual value. They constructed their dataset by extracting the three sentences most relevant to the environment from financial reports and associated them with the E score [^11] ESG scores are usually aggregated scores along multiple dimensions. In this study, they are focusing on the ”Environment” part of the ESG score.’s change and direction of change. Instead of relying on scores, clarkson_nlp_us_csr proposed to focus on good/bad CSR (corporate social responsibility) performers. Their approach evaluates CSR performance based on linguistic style rather than content, aiming to identify whether language patterns alone influence the perception of CSR quality. Greenscreen introduced a multi-modal dataset of the tweets of companies to predict a company’s ESG unmanaged risk. Finally, focusing on products and not companies, linSUSTAINABLESIGNALSijcai2023 introduced a dataset of online product descriptions and reviews used to predict the Finch Score. All these scores can be used to identify companies and products that are actually environmentally friendly, helping users distinguish actual sustainability from greenwashing. However, they can also be used to find discrepancies between the communicated green-ness of a product or company (e.g. the percentage of climate-related texts in reports) and the actual green-ness (e.g. ESG score or quantity of emissions), which might indicate cases of greenwashing.
Models
Mehra_2022 evaluated fine-tuned Transformers on ESG change prediction. clarkson_nlp_us_csr experimented with hand-chosen features plugged into random forest and SVM on CSR performers prediction. linSUSTAINABLESIGNALSijcai2023 also experimented with traditional approaches (e.g. gradient boosting) and custom architectures based on Transformer models on Finch score prediction. Greenscreen introduced baseline models, which are simple image and text embedding models (e.g., CLIP or sentence-BERT) with a classification head on ESG score prediction. bronzini_glitter_2023, clarkson_nlp_us_csr, Greenscreen did not report baselines, which makes the study difficult to analyze. Mehra_2022 reported good accuracy, demonstrating that a few sentences hold a significant amount of information to predict ESG score changes. However, there is definitely room for improvement, indicating that the information selected is not sufficient. lin-etal-2023-linear reported the performance of the baseline (average score) reaching already mean squared error (MSE) of 11.7% already quite low. Their Transformer-based approach reached MSE of 7.4% improving slightly on classical approaches (Gradient Boosting reaching MSE of 8.2%) [^12] Detailed performances can be found in Table 24 in the Appendix..
@@
We expected the prediction of ESG and CSR metrics to be difficult and require an extensive understanding of a company. However, bronzini_glitter_2023, linSUSTAINABLESIGNALSijcai2023, Mehra_2022 and clarkson_nlp_us_csr showed that it is possible to build strong predictors relying only on textual elements. This can be explained because analysts reward transparency [bronzini_glitter_2023], so the quantity and complexity of disclosure have strong predictive power [clarkson_nlp_us_csr]. Although existing studies have examined specific dimensions, future research should investigate the interaction between form [clarkson_nlp_us_csr] and content [bronzini_glitter_2023], as this could help uncover inconsistencies that may indicate greenwashing practices.
5 Greenwashing Detection
@@
There is a large body of work on disclosure (as described in Section 4.3.1). Quantifying climate-related disclosure can highlight greenhushing and selective disclosure. bingler_cheap_2021 identify a lack of disclosure in the Strategy and Metrics&Targets categories, which they describe as cherry-picking. They also highlight that merely announcing support for the TCFD does not lead to an increase in disclosure – on the contrary, it is a practice called cheap talk. Similarly, auzepy_evaluating_2023 conducted a more fine-grained analysis of TCFD-related disclosure in the banking industry. They also highlighted large differences across TCFD categories. In particular, they found a low disclosure rate on the fossil-fuel industry-related topic despite large investments in that sector.
@@
dingCarbonEmissionsTCFD2023 and chou_ESG found a correlation between climate-related disclosure and carbon emission: companies that emit more tend to disclose more climate-related information. bingler2023cheaptalkspecificitysentiment proposed a Cheap talk index based on claims specificity. They reached the same conclusion that larger emitters tend to disclose more, but also that negative news coverage is correlated with cheap talk. marco_polignano_nlp_2022 showed that while disclosing more, reports mostly focus on positive disclosure. hyewon_kang_analyzing_2022 proposed a sentiment ratio metric that highlights overly positive corporate reports. They identified periods of overly positive communications that followed negative environmental controversies; in other words, companies that tried to rebuild their image after a controversy. csr_report_greenwashing analyzed CSR reports of environmental violators and companies with clean records. They studied an ensemble of variables: quantification of environmental content, readability score, and sentiment analysis. They found that violators publish longer, more positive, and less readable CSR reports with more environment-related content. This suggests that companies use climate disclosure as a tool to mitigate controversies and/or to improve their image. kdir23 propose to use the discrepancies between internal disclosure and social media perception of a company to identify potential greenwashing.
@@
clarkson_nlp_us_csr found that companies that use a more complex language in their CSR disclosure are associated with a higher CSR rating. schimanski_bridging_2023 concluded that a higher quantity of ESG communication is associated with higher ESG rating. rouenEvolutionESGReports2023 also highlighted the relationship between disclosure quantity and complexity with the ESG score. This might indicate that the linguistic style is a good predictor of the substance of the discourse, or that analysts are rewarding companies that communicate on ESG-related issues.
@@
LEE_greenwashing proposed a greenwashing index based on the difference between the ESG score and the ESG score, weighted by the quantity of communication on each topic (E, S and G). Greenscreen define two types of risks: managed risk (the company is addressing it), and unmanaged risk (the company is not currently addressing it but it could). Based on these, they propose an ESG unmanaged risk score as a measure for environmental performance, which can then be used as a signal for identifying greenwashing.
@@
morio2023an proposed to evaluate a company’s stance towards climate change mitigation policies. They hypothesize that companies with a mixed stance might be engaging in greenwashing. The hypothesis was not tested and remains unconfirmed; still, such methodology helps understand the narratives of companies around climate change. coanComputerassistedClassificationContrarian2021 were able to analyze the type of contrarian claims used by conservative think-tank (CTTs) websites and contrarian blogs. They showed a shift over time from climate change denying arguments to opposing climate change mitigation policies. Despite this shift, they found that the primary donors of these think tanks continue to support organizations promoting a climate-denier narrative aimed at discrediting scientific evidence and scientists. Discrepancies between the stance extracted from official communications [morio2023an] on the one hand, and the stance of third-party media organizations such as think tanks founded by the company [coanComputerassistedClassificationContrarian2021] on the other, might indicate greenwashing.
@@
divinus_oppong-tawiah_corporate_2023 we propose to identify greenwashing using the linguistic profile of tweets. They estimate the deceptiveness of the text using a keyword-based approach and linguistic indicators (e.g. word quantity, sentence quantity). They found a correlation between greenwashing and lower financial performances. A scoring system based solely on linguistic elements has significant limitations. For example, the following tweet gets a high greenwashing score: “Read about #[company’s] commitment to a #lowcarbon future http://[company website]”. However, it is merely an invitation to read a Web page.
@@
Several indicators for greenwashing have been proposed: overly positive sentiment [hyewon_kang_analyzing_2022], the use of specific linguistic cues [divinus_oppong-tawiah_corporate_2023], discrepancies between ESG scores and corporate disclosures [LEE_greenwashing, Greenscreen], ambiguous stances [morio2023an], and inconsistencies between social media perception and official disclosures [kdir23]. While these approaches laid the theoretical groundwork for understanding indicators of greenwashing, they have a significant limitation: they are not empirically evaluated [measuring_greenwashing]. Only a few studies, such as that by hyewon_kang_analyzing_2022, have made attempts to validate their signals against actual cases, but even these efforts have not resulted in a comprehensive dataset of greenwashing examples. This gap between theory and practice highlights the necessity of developing datasets containing real-world instances of greenwashing. Without empirical evaluation, the prediction that a company engages in greenwashing is unfounded at best, and misleading or even defamatory at worst. However, building such a dataset comes with many challenges. First, one would have to identify suitable sources of documents that are openly accessible and likely to contain greenwashing. Second, one would have to overcome the inconspicuous and subjective nature of greenwashing itself. Finally, the publication of any such dataset exposes the authors to charges of defamation by the companies they accuse of greenwashing.
6 Open Challenges in Greenwashing Detection
While the lack of real-world greenwashing datasets constitutes a major obstacle, it is only one of several unresolved challenges. Building on the reviewed literature, this section examines additional limitations that affect greenwashing detection and, more broadly, NLP tasks characterized by ambiguity, subjectivity, and real-world constraints. Although discussed through the lens of greenwashing-related applications, these challenges are not specific to this domain and reflect broader methodological issues in applied NLP. We focus on three interconnected dimensions: (6.1) evaluation methodology, where ambiguity and task complexity complicate performance assessment; (6.2) model robustness, including sensitivity to noise and data imbalance; and (6.3) data sources, encompassing reproducibility, dataset availability, and definitional issues.
6.1 Evaluation Methodology
Firstly, as models improve, they are able to tackle more complex tasks, moving from classification to text generation, and from extraction to reasoning. This progression in task complexity creates an increase in evaluation complexity.
Classification Evaluation
Many tasks in greenwashing detection are framed as classification tasks. Such a framing may be too simplistic, in that it is blind to the semantic similarity between the classes. For example, in the Contrarian Claims dataset  [coanComputerassistedClassificationContrarian2021], models may fail on the fine-grained classification, but still get the super-category right. For such cases, morio2023an proposed strict and relaxed metrics, where the relaxed metric grouped similar labels together (e.g. ”strongly supporting” and ”supporting” grouped in ”supporting” label).